OpenStudy (anonymous):

solve for x 4000/x -10 = 4000/x+20

OpenStudy (igreen):

\(\dfrac{4000}{x} - 10 = \dfrac{4000}{x} + 20\) Is this it? @lyssaaaa

OpenStudy (anonymous):

no its\[\frac{ 4000 }{ x }-10=\frac{ 4000 }{ x+20 }\]

OpenStudy (igreen):

Oh, okay.

OpenStudy (igreen):

Okay x = -100 or 80.

OpenStudy (anonymous):

how did u get it @iGreen

OpenStudy (igreen):

You can give medals by clicking 'Best Response'. @lyssaaaa

OpenStudy (anonymous):

thank youuu